OpenStudy (anonymous):

What is the best first step to solve the system using elimination? x – 5y = 4 3x + 7y = –17 multiply the first equation by –3 multiply the second equation by 3 solve for x in the first equation solve for y in the second equation

OpenStudy (anonymous):

First of all you need to multiply the two y numbers so they can be the same like multiplying the first equation by 7 to get 7x-35y=28 and the second equation multiply it by 5 to get 15x+35y=-85. there you can cancel the y's to solve for x then substitute x for your answer to get y

OpenStudy (jkristia):

If those are your choises, then A; will give you 1: -3x + 15y = 12 2: 3x + 7y = -17 Now you can add the 2 and eliminate x, what you have is y = .... Then next step is to solve for x by replacing y with the value you just found